OpenStudy (anonymous):

Find the orthogonal trajectories of the family of curves. x^2 + 2y^2 = k^2

OpenStudy (anonymous):

first find dy/dx yes?

OpenStudy (anonymous):

im not sure if my dy/dx is correct.

OpenStudy (anonymous):

i got 

OpenStudy (anonymous):

that's right. So what is the slope that is perpendicular to that?

OpenStudy (anonymous):

its 2y/x

OpenStudy (anonymous):

so dy/dx = 2y/x yes?

OpenStudy (anonymous):

then i seperated the x and y's and got dy/2y = dx/x

OpenStudy (anonymous):

ok, integrate then what do you get?

OpenStudy (anonymous):

i got (1/2)ln(y) = 2(ln(x)) + C

OpenStudy (anonymous):

not sure if thats right haha.

OpenStudy (anonymous):

should be (1/2) lny = lnx + C

OpenStudy (anonymous):

oh yeah, i just simplified the y so i got 2(lnx + c)

OpenStudy (anonymous):

then y = Ce^(2lnx)

OpenStudy (anonymous):

but the book answer is different from this.

OpenStudy (anonymous):

simplify further

OpenStudy (anonymous):

e^2lnx = 2?

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

2LN X turns into lnx^2